# Prompt-Driven Building Footprint Extraction in Aerial Images with Offset-Building Model

### v3

@@ Abstract
More accurate extraction of invisible building footprints from very-high-resolution (VHR) aerial images relies on roof segmentation and roof-to-footprint offset extraction. Existing state-of-the-art methods based on instance segmentation suffer from poor generalization when extended to large-scale data production and fail to achieve low-cost human interactive annotation. The latest prompt paradigms inspire us to design a promptable framework for roof and offset extraction, which transforms end-to-end algorithms into promptable methods. Within this framework, we propose a novel Offset-Building Model (OBM). To rigorously evaluate the algorithm’s capabilities, we introduce a prompt-based evaluation method, where our model reduces offset errors by 16.6% and improves roof Intersection over Union (IoU) by 10.8% compared to other models. Leveraging the common patterns in predicting offsets, we propose Distance-NMS (DNMS) algorithms, enabling the model to further reduce offset vector loss by 6.5%. To further validate the generalization of models, we tested them using a new dataset with over 7,000 manually annotated instance samples. Our algorithms and dataset are available at https://anonymous.4open.science/r/OBM-B3EC.
@@ I Introduction
The problem of Building Footprint Extraction (BFE) has a history of over a decade, which can benefit 3D building modeling, building change detection, and building height estimation [1, 2, 3, 4, 5]. Early studies extract building footprints by measuring geometric features and traditional machine learning methods[6]. These kinds of methods[7],[8] usually limited by discriminating shallow features. Recently, deep learning based building related methods populate in solving BFE problems [9, 10, 11]. This kind of methods focus on near-nadir images where building roofs and footprints are nearly vertical or roof-to-footprint offsets are very small. However, demanding shotting angles of satellites make near-nadir images hard to be derived.
@@
To solve this problem, Wang et al. [12] propose a model called Learning Offset Vector (LOFT) and a related dataset BONAI in 2023. LOFT provides a new idea to solve the BFE problem. It simulates footprint-labelling process of human, extracting a roof and a roof-to-footprint offset, and then utilizing both outputs, directly drag the roof to its footprint in the direction of offset. The LOFT model adopts a two-stage instance segmentation structure similar to Mask RCNN[13], featuring an offset head based on convolution and Region of Interest Align (ROI Align). Based on his idea, BFE problem found a new solution.
@@
However, when extending such method to large-scale data production, such idea suffers from poor generalization. As illustrated in Fig. 1, this image is from real data production, whose shooting location and other camera related information are unknown towards the model. Apart from plenty of mistakes made by instance-level models, hyper-parameters of NMS often let data producers in a dilemma: higher output score thresholds of NMS will miss more samples, while lower scores let them hard to select correct instances.
Refer to caption LOFT + NMS
Refer to caption LOFT + soft NMS
Refer to caption LOFT + soft NMS + Merge
Refer to caption Ground Truth
@@
Figure 1: In given pictures, red boundaries and green boundaries represent roof and footprint respectively. During large-scale data production, instance segmentation methods face challenges related to generalization. The listed results are from LOFT[12], and the input images are from real production process which is 100% unseen by LOFT. Apart from mistake recognition, problems manifest in two aspects. Firstly, these methods usually rely on postprocessing algorithms. Showing in first picture, a strict NMS algorithm lost many instances. To address this, soft NMS[14] often applied to minimize the number of missing samples. However, lower score thresholds of soft NMS consequently matched one building with many instances in pieces. The confusing results let data producer hard to choose correct instances. Of course, predicted buildings in neighbor can be merged and fused together as shown in the third picture. However, the results let densely populated buildings stick together, far from getting wanted results as listed ground truth. Secondly, data producers have to plot out those missing samples point-by-point, because of inflexible Region Proposal Network (RPN).
@@
With the turnup of prompt paradigm, a cutting-edge model, namely Segment Anything Model (SAM)[16], allows for powerful zero-shot capabilities with flexible prompting. SAM has applied widely in dataset labeling. Data producers only plot a point, a box or a coarse-labeled mask, then SAM can give a better mask label. Inherit from which, many works have proved its success. Finetuned from SAM, SAM-HQ improved segmentation quality of SAM by solving the problem of 16$\times$-downsampling feature maps[17]. RS-Prompter[18] can complete many downstream mask-related tasks by providing prompts to SAM. Integrating how Wang[12] finds building footprints and powerful SAM, we are lighted to solve BFE problem in prompt level: “can we provide a prompt and a model directly plots the roof and drag it to its footprint like a human annotator?”.
To answer mentioned two questions, we devised a workflow inspired by SAM, transforming Wang’s models[12] into ones compatible with box prompts. Additionally, leveraging SAM, we developed our own model, Offset Building Model (OBM). To tackle problems of NMS algorithms, we identified general patterns by experiments, leading to the creation of the DNMS algorithms. DNMS will not delete instances, but improve offsets quality in reality. We also established a novel evaluation system for prompt-level models, by which we can depict ability of models in more details. By conducting extra experiments on a new dataset, we proved the superiority of our OBM and DNMS algorithms.
@@
Moreover, with accurate extraction of roofs and footprints, another contribution of our work can influence the process of monocular height estimation. Height information can be retrieved directly with 3D-aware techniques, e.g., 3D sensors such as light detection and ranging (LiDAR)[19, 20, 21]. However, rare data derived by such conditionally applicable sensors cannot fulfill the demand of related deep learning methods[5], but our methods can generate relative height map of buildings using optical remote sensing images. Different from above mentioned height maps, the only gap between height maps and our maps can bridged by a scale factor or camera parameters.
In summary, our contribution is as follows:
@@ 1. $\bullet$
    We proposed a framework, under which can using prompts to extract building footprint by predicting an offset and a roof segmentation. Under this framework, we also propose a powerful model, named OBM, as Fig. 2.
@@ 2. $\bullet$
    We designed DNMS algorithms to improve the quality of predicted offsets.
@@ 3. $\bullet$
    We presented an effective and detailed metric method for prompt-level offset and roof regression.
@@ II Related Work
OBM can be seen as an application of a large vision model (SAM) and multi-task learning to solve BFE problem. Similar with object query, OBM uses an offset query to represent roof-to-footprint offset.
@@
Segment Anything Model and Applications: Segmentation is a fundamental task in computer vision[22, 23, 24, 25, 26, 27, 28, 16], which requires a pixel-wise understanding of a image. SAM is a powerful zero-shot segmentation model[16], by using minimal human input, such as bounding boxes, reference points, or simply text-based prompts. Based on SAM, Fast-SAM[25] realized a 50x speedup by using a CNN-based decoder. PerSAM can segment personal objects in any context with favorable performance, and PerSAM-F further alleviates the ambiguity issue by scale-aware fine-tuning[29]. PerSAM can also improve DreamBooth[30] by mitigating the disturbance of backgrounds during training. SAM-PT[31] prompts SAM with sparse point trajectories predicted using point trackers, such as CoTracker[32]. SAM-HQ[17] is a high-quality version of SAM, which can generate high-resolution masks. This improvement is achieved by introducing negligible overhead to the original SAM. In SAM-HQ, a lightweight High-quality Output Token in HQ-SAM to replace the original SAM’s output token for high-quality mask prediction. SAM-Adapter[33] proposes the adaptation approach to adapt SAM to downstream tasks and achieve enhanced performance. After SAM launched, it has been used in remote sensing [34, 18]. In [34], involves using a prompt-text-based segmentation as a training sample (instead of a human-labeled sample), making it an automated process for refining SAM on remote sensing imagery. RSprompter[18] proposes a novel prompt learning method that augments the SAM model’s capabilities, thereby facilitating instance segmentation in remote sensing imagery. SAM has not be applied in building footprint extraction, and OBM is the first work to apply SAM in BFE problem.
@@
BFE Problem Solutions: Spatial residual inception (SRI) module[35] was proposed to capture and aggregate multi-scale contextual information, and this module could improve the discriminative ability of the model and obtain more accurate building boundaries. Jian Kang[36] investigates the problem of deep learning-based building footprint segmentation with missing annotations, approaching it from the perspective of designing an effective loss function to specifically deal with this issue. Weijia Li[37] explored the combination of multi-source GIS map datasets and multi-spectral satellite images for building footprint extraction in four cities using a U-Net-based semantic segmentation model for building footprint extraction. ConvBNet[38] and MHA-Net[39] was tailored to the complex textures, varying scales and shapes, and other confusing artificial objects in building footprint extraction. Apart from that, building segmentation is also a related task[40, 41, 42, 43]. DARNet[40] using proposed loss function to directly encourages the contours to match building boundaries. Zhanming Ma et al.[41] proposed a local feature search network with discarding attention module (DFSNet) to help the model distinguish the building areas and water bodies. YileiShi et al.[42] integrating graph convolutional network (GCN) and deep structured feature embedding (DSFE) into an end-to-end workflow to overcome the issue of delineation of boundaries. RIU-Net[43], a residually connected, Inception-based, u-shaped encoder-decoder architecture with skip connections, was proposed to segment buildings in remote sensing imagery. Those semantic methods can not predict a fine-grained building footprint for those high-rise buildings, while LOFT exists with many problems as mentioned in Section I.
@@
Object Query: Queries, stored as learnable vectors, are widely used to imply properties of instance[44, 45, 46, 47, 48, 49, 50]. Object query for detection was first proposed in DETR[45], which is used to represent the bounding box area. In DETR, the model will pre-set $n\times$ object queries for $n$ instances, and, at the finally stage, these $n$ queries will be decoded as $n$ bounding boxes with several layers of Feed-Forward Network (FFN). Followed by that, MaskFormer[46] and Mask2Former[47] are another series of transformer model for semantic segmentation, using queries to represent instances. Both models use a single head to predict binary masks and a classification head to predict the class of each mask layer. Dynamic Anchor Boxes DETR(DAB-DETR)[44] offers a deeper understanding of the role of queries in DETR. DeNoising DETR(DN-DETR)[48] offered a deepened understanding of the slow convergence issue of DETR-like methods, and using a novel denoising method related with queries to speedup training. Based on aforementioned ideas, the same team proposed DINO[49] for object detection and MaskDINO[50] for instance segmentation, semantic segmentation and panoptic segmentation. OBM is the first work to use offset query and ROAM structure to predict roof-to-footprint offset.
@@
Multi-Task Learning: Multi-Task Learning (MTL) is a learning paradigm in machine learning and its aim is to leverage useful information contained in multiple related tasks to help improve the generalization performance of all the tasks [51]. Multi-task model can do more things using the same model [52, 12, 53, 54, 55]. Hybrid Task Cascade (HTC)[52] effectively integrates cascade into instance segmentation by interweaving detection and segmentation features together for a joint multi-stage processing. LOFT[12] can predict roof-to-footprint offset with roof segmentation. P. Li et al.[53] proposed a method which can extract road and road intersections. Eagle-Eyed Multitask CNNs[54] incorporate three tasks, aerial scene classification (ASC), center-metric learning and similarity distribution learning together. For autonomous driving, a shared backbone with different subnets was designed to address object detection, drivable area segmentation, lane detection, depth estimation and so on [55]. HybridNet [56] is one of those models developed an end-to-end perception network to perform multi-tasking, including traffic object detection, drivable area segmentation and lane detection. Additionally, YOLOP [57] and YOLOPv2 [58] further enhanced the performance of aforementioned three tasks by using the share-weight backbone.
@@ III-B Offset Building Model (OBM)
OBM can predict roof-to-footprint offsets, roof segmentation, and building segmentation for each inputted image and its bounding box prompts. Using predicted offsets and roofs, a relative height map and building footprints can be figure out.
@@
As shown in Fig. 3, OBM inherits from SAM’s segmentation ability and we implement a novel Reference Offset Augment Module (ROAM) to predict roof-to-footprint offsets. ROAM includes a series of offset tokens, FFN and offset coders. To match different capability of GPUs and accelerate training, we designed a Prompt Sampler. In OBM, we preserve most of SAM structures for its ability of perceiving objects. Since the roof and building have sharing semantic pixels, what we need to do is ensuring that each mask token focuses solely on the roof or the building. In other words, eliminating cognitive ambiguity in models.

### v4

@@ Abstract
More accurate extraction of invisible building footprints from very-high-resolution (VHR) aerial images relies on roof segmentation and roof-to-footprint offset extraction. Existing methods based on instance segmentation suffer from poor generalization when extended to large-scale data production and fail to achieve low-cost human interaction. This prompt paradigm inspires us to design a promptable framework for roof and offset extraction, and transforms end-to-end algorithms into promptable methods. Within this framework, we propose a novel Offset-Building Model (OBM). Based on prompt prediction, we first discover a common pattern of predicting offsets and tailored Distance-NMS (DNMS) algorithms for offset optimization. To rigorously evaluate the algorithm’s capabilities, we introduce a prompt-based evaluation method, where our model reduces offset errors by 16.6% and improves roof Intersection over Union (IoU) by 10.8% compared to other models. Leveraging the common patterns in predicting offsets, DNMS algorithms enable models to further reduce offset vector loss by 6.5%. To further validate the generalization of models, we tested them using a newly proposed test set, Huizhou test set, with over 7,000 manually annotated instance samples. Our algorithms and dataset will be available at https://github.com/likaiucas/OBM.
@@ I Introduction
The problem of Building Footprint Extraction (BFE) has a history of over a decade, which can benefit 3D building modeling, building change detection, and building height estimation [1, 2, 3, 4, 5], and there were many meaningful researches about cities[6, 7, 8]. Early studies extract building footprints by measuring geometric features and traditional machine learning methods[9]. These kinds of methods[10, 11] usually limited by discriminating shallow features and cannot be applied in a larger scale. Recently, deep learning based building related methods populate in solving BFE problems [12, 13, 14]. This kind of methods focus on near-nadir images where building roofs and footprints are nearly vertical or roof-to-footprint offsets are very small. However, demanding shooting angles of satellites make near-nadir images hard to be derived. In other words, if models can extract footprints in off-nadir images, the production cost of obtaining remote sensing images through photography will be lowered.
@@
To solve this problem, Wang et al. [15] propose a model called Learning Offset Vector (LOFT) and a related dataset BONAI in 2023. LOFT provides a new idea to solve the BFE problem. It simulates footprint-labelling process of human, extracting a roof and a roof-to-footprint offset, and then utilizing both outputs, directly drag the roof to its footprint in the direction of offset. The LOFT model adopts a two-stage instance segmentation structure similar to Mask RCNN[16], featuring an offset head based on convolution and Region of Interest Align (ROI Align). Under the similar idea of LOFT, Weijia Li proposed MLS-BRN[17], a multi-level supervised building reconstruction network which allows more kinds of annotated building dataset involved in training the same model.
@@
However, when extending such method to large-scale data production, such idea suffers from poor generalization. Additionally, the unstable performance of NMS algorithms will make the outputs of these end-to-end models hardly be applied in data production. As illustrated in Fig. 1, this image is from real data production, whose shooting location and other camera related information are unknown towards the model. Within the given images, apart from plenty of mistakes made by instance-level models, hyperparameters of NMS often let data producers in a dilemma: higher output score thresholds of NMS will miss more samples, while lower scores let them hard to select correct instances.
LOFT + NMS
LOFT + soft NMS
LOFT + soft NMS + Merge
Ground Truth
@@
Figure 1: In given pictures, red boundaries and green boundaries represent roof and footprint respectively. During large-scale data production, instance segmentation methods face challenges related to generalization. The listed results are from LOFT[15], and the input images are from real production process which is 100% unseen by LOFT. Apart from mistake recognition, problems manifest in two aspects. Firstly, these methods usually rely on post-processing algorithms. Showing in first picture, a strict NMS algorithm lost many instances. To address this, soft NMS[18] often applied to minimize the number of missing samples. However, lower score thresholds of soft NMS consequently matched one building with many instances in pieces. The confusing results let data producer hard to choose correct instances. Of course, predicted buildings in neighbor can be merged and fused together as shown in the third picture. However, the results let densely populated buildings stick together, far from getting wanted results as listed ground truth. Secondly, data producers have to plot out those missing samples point-by-point, because of inflexible Region Proposal Network (RPN).
With the turn up of prompt paradigm, a cutting-edge model, namely Segment Anything Model (SAM)[19], allows for powerful zero-shot capabilities with flexible prompting. SAM has applied widely in dataset labeling. Data producers only plot a point, a box or a coarse-labeled mask, then SAM can give a better mask label. Inherit from which, many works have proved its success. Fine-tuned from SAM, SAM-HQ improved segmentation quality of SAM by solving the problem of 16$\times$-downsampling feature maps[20]. RS-Prompter[21] can complete many downstream mask-related tasks by providing prompts to SAM. If we can control the input of prompts and the model will only give results in interested areas, the dilemma caused by end-to-end methods will be solved. Integrating how Wang[15] finds building footprints and powerful SAM, we are lighted to solve BFE problem in prompt level: “can we provide a prompt and a model directly plots the roof and drag it to its footprint like a human annotator?”.
@@
To answer mentioned two questions, we devised a workflow inspired by SAM, transforming Wang’s models[15] into a promptable model compatible with box prompts. Additionally, leveraging SAM, we proposed Offset Building Model (OBM)Fig. 2. To tackle problems of NMS algorithms, we identified general patterns by experiments, leading to the creation of the DNMS algorithms. DNMS will not delete instances, but improve offsets quality in reality. We also established a novel evaluation system for prompt-level models, by which we can depict ability of models in more details. By conducting extra experiments on a new dataset, we proved the superiority of our OBM and DNMS algorithms.
@@
Moreover, with accurate extraction of roofs and footprints, another contribution of our work can influence the process of monocular height estimation. Height information can be retrieved directly with 3D-aware techniques, e.g., 3D sensors such as light detection and ranging (LiDAR)[23, 24, 25]. However, rare data derived by such conditionally applicable sensors cannot fulfill the demand of related deep learning methods[5], but our methods can generate relative height map of buildings using optical remote sensing images. Different from above-mentioned height maps, the only gap between height maps and our maps can be bridged by a scale factor or camera parameters.
In summary, our contributions are as follows:
@@ 1. $\bullet$
    We proposed a prompt-level footprint model OBM, and OBM was the first model which introduced the offset tokens and ROAM structure to predict offsets.
@@ 2. $\bullet$
    Based on the idea of prompt, we transformed open-source RoI-based methods into prompt mode.
@@ 3. $\bullet$
    We designed DNMS algorithms to improve the quality of predicted offsets. The algorithms are based on the common pattern of predicting offsets and can effectively improve the directional quality of offsets.
  4. $\bullet$
    We presented a comprehensible and detailed metric method for prompt-level offset and roof regression.
@@ II Related Work
OBM can be seen as an application of a large vision model (SAM) and multitask learning to solve BFE problem. Similar with object query, OBM uses an offset query to represent roof-to-footprint offset.
@@
Segment Anything Model and Applications: Segmentation is a fundamental task in computer vision[26, 27, 28, 29, 30, 31, 32, 19], which requires a pixel-wise understanding of a image. SAM is a powerful zero-shot segmentation model[19], by using minimal human input, such as bounding boxes, reference points, or simply text-based prompts. Based on SAM, Fast-SAM[29] realized a 50x speedup by using a CNN-based decoder. PerSAM can segment personal objects in any context with favorable performance, and PerSAM-F further alleviates the ambiguity issue by scale-aware fine-tuning[33]. PerSAM can also improve DreamBooth[34] by mitigating the disturbance of backgrounds during training. SAM-PT[35] prompts SAM with sparse point trajectories predicted using point trackers, such as CoTracker[36]. SAM-HQ[20] is a high-quality version of SAM, which can generate high-resolution masks. This improvement is achieved by introducing negligible overhead to the original SAM. In SAM-HQ, a lightweight High-quality Output Token in HQ-SAM to replace the original SAM’s output token for high-quality mask prediction. SAM-Adapter[37] proposes the adaptation approach to adapt SAM to downstream tasks and achieve enhanced performance. After SAM launched, it has been used in remote sensing [38, 21]. In [38], involves using a prompt-text-based segmentation as a training sample (instead of a human-labeled sample), making it an automated process for refining SAM on remote sensing imagery. RSprompter[21] proposes a novel prompt learning method that augments the SAM model’s capabilities, thereby facilitating instance segmentation in remote sensing imagery. SAM-RBox can detect a rotated box for different items in remote sensing images, which uses a trained horizontal Fully Convolutional One-Stage object (FCOS) detector to provide HBoxes into SAM as prompts [39]. SAM has not be applied in building footprint extraction, and OBM is the first work to apply SAM in BFE problem.
@@
BFE Problem Solutions: Spatial residual inception (SRI) module[40] was proposed to capture and aggregate multiscale contextual information, and this module could improve the discriminative ability of the model and obtain more accurate building boundaries. Jian Kang[41] investigates the problem of deep learning-based building footprint segmentation with missing annotations, approaching it from the perspective of designing an effective loss function to specifically deal with this issue. Weijia Li[42] explored the combination of multi-source GIS map datasets and multi-spectral satellite images for building footprint extraction in four cities using a U-Net-based semantic segmentation model for building footprint extraction. ConvBNet[43] and MHA-Net[44] was tailored to the complex textures, varying scales and shapes, and other confusing artificial objects in building footprint extraction. Apart from that, building segmentation is also a related task[45, 46, 47, 48]. DARNet[45] using proposed loss function to directly encourages the contours to match building boundaries. Zhanming Ma et al.[46] proposed a local feature search network with discarding attention module (DFSNet) to help the model distinguish the building areas and water bodies. Yilei Shi et al.[47] integrating graph convolutional network (GCN) and deep structured feature embedding (DSFE) into an end-to-end workflow to overcome the issue of delineation of boundaries. RIU-Net [48], a residual connected, Inception-based, u-shaped encoder-decoder architecture with skip connections, was proposed to segment buildings in remote sensing imagery. In offset-based methods, Gordon Christie[49] introduced a U-Net based methods to predict image-level orientation as $sin(\theta)$ and $cos(\theta)$. With this map, a given ground truth footprint or roof can finally be dragged to its related roof or footprint. Using a similar idea, MTBR-Net[50] using semantic related tasks and offset related tasks to extract 3D buildings and was equipped with more functions. Based on a similar idea of offsets, LOFT[15] uses instance-level offset as a feature to describe buildings. In CVPR 2024, MLS-BRN[17] adds new tasks to bridge the gaps between different building instance, which finally alleviates the demand on 3D annotations. Different with aforementioned, OBM supports a prompt-level footprint extraction and using DNMS algorithms to improve the prediction quality of offsets.
@@
Object Query: Queries, stored as learnable vectors, are widely used to imply properties of instance[51, 52, 53, 54, 55, 56, 57]. Object query for detection was first proposed in DETR[52], which is used to represent the bounding box area. In DETR, the model will pre-set $n\times$ object queries for $n$ instances, and, at the final stage, these $n$ queries will be decoded as $n$ bounding boxes with several layers of Feed-Forward Network (FFN). Followed by that, MaskFormer[53] and Mask2Former[54] are another series of transformer model for semantic segmentation, using queries to represent instances. Both models use a single head to predict binary masks and a classification head to predict the class of each mask layer. Dynamic Anchor Boxes DETR(DAB-DETR)[51] offers a deeper understanding of the role of queries in DETR. DeNoising DETR(DN-DETR)[55] offered a deepened understanding of the slow convergence issue of DETR-like methods, and using a novel denoising method related with queries to speedup training. Based on aforementioned ideas, the same team proposed DINO[56] for object detection and MaskDINO[57] for instance segmentation, semantic segmentation and panoptic segmentation. OBM is the first work to use offset query and Reference Offset Augment Module (ROAM) structure to predict roof-to-footprint offset.
@@
Multi-Task Learning and Deep Learning in Remote Sensing: Multi-Task Learning (MTL) is a learning paradigm in machine learning and its aim is to leverage useful information contained in multiple related tasks to help improve the generalization performance of all the tasks [58]. Multitask model can do more things using the same model [59, 15, 60, 61, 62, 63]. Hybrid Task Cascade (HTC)[59] effectively integrates cascade into instance segmentation by interweaving detection and segmentation features together for a joint multi-stage processing. LOFT[15] can predict roof-to-footprint offset with roof segmentation. P. Li et al.[60] proposed a method which can extract road and road intersections. Eagle-Eyed Multitask CNNs[61] incorporate three tasks, aerial scene classification (ASC), center-metric learning and similarity distribution learning together. For autonomous driving, a shared backbone with different subnets was designed to address object detection, drivable area segmentation, lane detection, depth estimation and so on [62]. HybridNet [64] is one of those models developed an end-to-end perception network to perform multi-tasking, including traffic object detection, drivable area segmentation and lane detection. Additionally, YOLOP [65] and YOLOPv2 [66] further enhanced the performance of aforementioned three tasks by using the share-weight backbone. ABNet[63] was designed to solve challengings caused by complicated backgrounds and imbalanced scale and distribution of remote sensing image. Salient object detection in optical remote sensing images often suffers from intrinsic problems, such as cluttered background, scale variation and irregular topology, to solve this HFANet[67] was proposed. SDNet[68] is a dual-branch network which can perform cross-task knowledge distillation from the scene classification to facilitate accurate saliency detection.
@@ III-B Offset Building Model (OBM)
OBM can predict roof-to-footprint offsets, roof segmentation, and building segmentation for each inputted image and its bounding box prompts. Using predicted offsets and roofs, a relative height map and building footprints can be figured out.
@@
As shown in Fig. 3, OBM inherits from SAM’s segmentation ability, and we implement a novel Reference Offset Augment Module (ROAM) to predict roof-to-footprint offsets. ROAM includes a series of offset tokens, FFN and offset coders. To match different capability of GPUs and accelerate training, we designed a Prompt Sampler. In OBM, we preserve most of SAM structures for its ability of perceiving objects. Since the roof and building have sharing semantic pixels, what we need to do is ensuring that each mask token focuses solely on the roof or the building. In other words, eliminating cognitive ambiguity in models.